# Harmonizing SO(3)-Equivariance with Neural Expressiveness: a Hybrid Deep Learning Framework Oriented to the Prediction of Electronic Structure Hamiltonian

### v9

@@ Abstract
Deep learning for predicting the electronic structure Hamiltonian of quantum systems necessitates satisfying the covariance laws, among which achieving SO(3)-equivariance without sacrificing the non-linear expressive capability of networks remains unsolved. To navigate the harmonization between equivariance and expressiveness, we propose a deep learning method, namely HarmoSE, synergizing two distinct categories of neural mechanisms as a two-stage cascaded regression framework. The first stage corresponds to group theory-based neural mechanisms with inherent SO(3)-equivariant properties prior to the parameter learning process, while the second stage is characterized by a non-linear 3D graph Transformer network we propose featuring high capability on non-linear expressiveness. The novel combination lies in the point that, the first stage predicts baseline Hamiltonians with abundant SO(3)-equivariant features extracted, assisting the second stage in empirical learning of equivariance; and in turn, the second stage refines the first stage’s output as a fine-grained prediction of Hamiltonians using powerful non-linear neural mappings, compensating for the intrinsic weakness on non-linear expressiveness capability of mechanisms in the first stage. Our method enables precise, generalizable predictions while maintaining robust SO(3)-equivariance under rotational transformations, and achieves state-of-the-art performance in Hamiltonian prediction on six benchmark databases.
keywords:
Deep Learning, Neural Networks, AI for Science, Atomic Simulation, 3D Graphic Modeling
1 Introduction
@@
Recently, deep learning methods [19, 21, 7, 13, 27, 6] have emerged as a promising trend for predicting the electronic structure Hamiltonian, an essential physical quantity in understanding a wide range of properties, including electronic structures, magnetic properties, optics, transport, and numerous other properties. These methods have demonstrated great potentials in terms of prediction accuracy and offer a way to bypass the computationally exhaustive self-consistent steps of the traditional Density Functional Theory (DFT) method [9, 11], thereby providing a viable pathway for the effective simulation and design of large-scale atomic systems, laying the foundation for many down-stream applications [24], particularly in the development of hardware devices that are critical for information processing, storage, and computation.
@@
However, applying deep learning techniques to the Hamiltonian prediction task continues to present substantial challenges, highlighting the evolving demands on deep learning technology. An ultra-high accuracy to $10^{-1}$ meV scale ($10^{-4}$ eV) is required to precisely calculate the down-stream physical quantities such as band structures. Furthermore, the fidelity of Hamiltonian predictions should not be confined to a specific coordinate system; rather, the results must demonstrate robust symmetry and generalizability across various choices of reference frames. This necessitates that deep learning methods capture the intrinsic symmetries of the Hamiltonian with respect to transformations of the coordinate system, ensuring the consistency with physical principles covariant to these transformations. In the context of Hamiltonian prediction, the key covariance principles are 3D translational invariance and rotational equivariance. By utilizing relative coordinates, 3D translational invariance has been effortlessly achieved, while 3D rotational equivariance, i.e., equivariance to the SO(3) group, remains a challenging target to guarantee. This difficulty arises because the Hamiltonian of each pair of atoms is usually high-dimensional, and its variation space under rotational disturbance is large. Consequently, it is difficult to cover the vast variability space they inhabit merely depending on parameter learning from discrete training samples. To address this, several works [27, 6] applied group theory-guaranteed feature descriptors and tensor operators assuring inherent SO(3)-equivariance prior to the data-driven parameter learning process. However, to guarantee such SO(3)-equivariance independent to specific network parameters, these methods highly restricted the use of non-linear activation layers for SO(3)-equivariant features, i.e. those with degrees greater than 00. Given that the expressiveness capability of modern deep networks heavily relies on the extensive use of non-linear mappings, this restriction resulted in a significant drawback on network’s expressiveness capabilities. This, in turn, led to bottlenecks in fitting complex mappings, limiting the accuracy achievable in predicting Hamiltonians. This dilemma, is also broadly prevalent in other 3D machine learning tasks where SO(3)-equivariance is highlighted, as analyzed by Zitnick et al. [28].
@@
To harmonize SO(3)-equivariance and expressiveness for the prediction of electronic structure Hamiltonians, this paper proposes a novel two-stage regression framework, namely HarmoSE, which combines mechanisms boasting parameter-indep-endent, prior SO(3)-equivariance with mechanisms featuring flexibility in non-linear expressiveness, overcoming the respective challenges of each categories of mechanisms, i.e., the limited non-linear expressive capability for the former as well as the difficulty of learning SO(3)-equivariance from data for the latter, through effective complementary strategies. Specifically, the first regression stage corresponds to the neural mechanisms constructed based on group theory with prior equivariant properties of 3D atomic systems, predicting an approximate value of the Hamiltonian, with abundant SO(3)-equivariant features provided. In the second stage, a highly expressive graph Transformer network we design, with no restrictions on non-linear activations, takes over. This network dynamically learns the 3D structural patterns of the atomic systems, compensates for the expressiveness shortcomings of the first-stage network arising from limited non-linear mappings, and refines the Hamiltonian values predicted in the first stage to enhance accuracy. Although this stage might not possess a parameter-independent prior SO(3)-equivariance due to its non-linearity, it is capable of capturing SO(3)-equivariance through learning effective network parameters via the help of three pivotal mechanisms. First, instead of directly regressing the entire Hamiltonians, the second stage aims to refine the Hamiltonian predictions from the first stage with corrective adjustments in a cascaded manner. The scopes of adjustments are smaller, lowing down the difficulties on non-linear learning of SO(3)-equivariance. Second, the second-stage network incorporates covariant features, including SO(3)-equivariant features extracted by the first-stage network and SO(3)-invariant features engineered by geometric knowledge, with its inputs to assist in the implicit learning of SO(3)-equivariance. Third, as the core of Transformer, the attention mechanism has potentials to adapt to geometric condition variations such as coordinate transformations, through its dynamic weighting strategy. Collectively, the combination of the two categories of neural mechanisms in the two stages allows the framework to overcome the challenges of each individual mechanism and make precise, generalizable, and SO(3)-equivariant predictions, being much more effective than simply increasing the parameter count for the network from one stage and fine-tuning it alone.
Our contributions are summarized below:
• Our work explores new deep learning paradigm with the synergy of two categories of distinct neural mechanisms, where the first one exhibits prior equivariance independent of parameter learning, and the second one possesses ample non-linear expressiveness, overcoming the respective challenges each mechanism faces via fully releasing their complementarity, and successfully harmonizing SO(3)-equivariance with neural expressiveness.
•  We design a Transformer architecture within the proposed framework, incorporating advanced mechanisms such as covariant feature integration and multi-head attentions, enabling effective modeling of diverse 3D atomic structures with SO(3)-equivariant properties.
• Our method achieves state-of-the-art (SOTA) performance on Hamiltonian regression in six benchmark databases, showing good potential on simulating diverse categories of atomic systems. Particularly, our SOTA performance in the twisted samples, which exhibit SO(3)-equivariance effects due to the inter-layer rotations, comprehensively confirm the strong capability in capturing the intrinsic SO(3)-equivariance of Hamiltonians. The materials consisted in these databases have extensive values in the realm of information technology, implicating an applicable potential of our method in advancing the development and optimization of information hardware.
2 Related Work
@@
As representative researches on equivariance to discrete rotational group, Dieleman et al. [3] introduced cyclic symmetry operations into CNNs to achieve rotational equivariance; Ravanbakhsh et al. [18] explored parameter-sharing techniques for equivariance to discrete rotations; Kondor et al. [12] developed equivariant representations via compositional methods and tensor theory; Zitnick et al. [28], Passaro et al. [17] and Liao et al. [15] introduced spherical harmonic bases for atomic modeling, focusing on rotational equivariance but limited to discrete sub-groups of SO(3) due to their sampling strategy. These approaches were very effective on discrete symmetries but sub-optimal on handling continuous 3D rotations. Focusing on equivariance to continuous rotational group, Jaderberg et al. [10] and Cohen et al. [2] achieved considerable success in 2D image recognition tasks by modeling equivariance to in-plane rotations. However, their applications were limited within the scope of 2D tasks and did not fit for the more complex demands of equivariance to 3D continuous rotational group, i.e. SO(3), required in the Hamiltonian prediction task.
@@
In the field of researches on equivariance to SO(3), approaches like DeepH [13] explored equivariance via a local coordinate strategy, which made inference within the fixed local coordinate systems built with neighboring atoms, then transferred the output according to equivariance rules to the corresponding global coordinates. However, due to a lack of in-depth exploration of SO(3)-equivariance at the neural mechanism level, this method faced challenges when the local coordinate system underwent rotational disturbances from non-rigid deformation, e.g. the inter-layer twist of bilayer structures. In contrast, some methods [20, 4, 5] considered SO(3)-equivariance from the perspective of intrinsic mechanisms of neural networks, developed group theory-informed equivariant operations, such as linear scaling, element-wise sum, direct sum, direct product, as well as the Clebsch-Gordan decomposition, effectively applying in tasks like molecular dynamics simulation [16, 14] and Hamiltonian prediction [19, 21, 6], where DeepHE3 [6] was validated as a SOTA method across diverse atomic systems. However, a common challenge across these methods lies in the fact that, to achieve inherent equivariance prior to the parameter learning process, they forbade the use of complex non-linear mappings like $Sigmoid$, $SiLU$, and $Softmax$ for SO(3)-equivariant features, significantly limiting the network’s expressive potential and creating a bottleneck in generalization performance. Although these approaches tried to enhance expressiveness via a gated activation function, where SO(3)-invariant features undergone through non-linear activation layers were used as gating coefficients that were multiplied with SO(3)-equivariant features, this mechanism, viewed from the perspective of equivariant features, amounted to a linear operation and did not fundamentally improve their expressive capability. For these methods, this equivariance-expressiveness dilemma remains an unsolved problem.
3 Preliminary
@@
In the study of symmetry on mathematical structures, covariance encompasses two fundamental concepts: invariance and equivariance. An operation $A$ is called invariant under another operation $B$ if the result of $A$ remains unchanged when $B$ is applied, formally $A(B(x))=A(x)$; on the other hand, $A$ is equivariant with respect to $B$ if applying $B$ before or after $A$ has the same effect, expressed as: $A(B(x))=B(A(x))$. The key covariance properties of Hamiltonians are the 3D translational invariance and rotational equivariance with respect to reference frame. Since 3D translational invariance can be easily achieved by using the relative coordinates between two atoms, the focus here should be on studying equivariance to SO(3), i.e. a group consisting of all continuous 3D rotation operations. In the context of Hamiltonians, when the reference frame rotates by a rotation matrix denoted as $\mathbf{R}$, the edge of an atom pairs $(i,j)$ transforms from $\mathbf{r}_{ij}$ to $\mathbf{R}\cdot\mathbf{r}_{ij}$, and the Hamiltonians in the direct sum state transforms equivariantly from $\mathbf{h}_{ij}$ to $D(\mathbf{R})\cdot\mathbf{h}_{ij}$, where $D(\mathbf{R})$ is the Wigner-D matrix. Note that here we present SO(3)-equivariance under the direct sum state due to its simple vector form. For the equivalent formulation under the matrix-formed direct product state, please refer to Gong et al. [6].
The requirements for the fitting and generalization capability of a neural network $f_{nn}(\cdot)$ for Hamiltonian prediction can be formally expressed as: $f_{nn}(\{\mathbf{r}_{ij}|i\in Nodes,ij\in Edges\})\cong\{\mathbf{h}_{ij}\}$; moreover, the requirement on SO(3)-equivariance can be represented as: $f_{nn}(\{\mathbf{R}\cdot\mathbf{r}_{ij}|i\in Nodes,ij\in Edges\})\cong\{D(\mathbf{R})\cdot\mathbf{h}_{ij}\}$. It is crucial that $f_{nn}(\cdot)$ intrinsically embodies SO(3)-equivariance to effectively generalize under rotational reference frames. An intuitive way to achieve this is data-driven empirical learning, however, this is challenging for the Hamiltonian prediction task, since $\mathbf{h}_{ij}$, consisting of multiple angular momentums, is typically high-dimensional with a total of dimensions likely reaching into the hundreds or thousands. Moreover, $\mathbf{h}_{ij}$ is not determined in isolation by the types of atom $i$ and $j$ as well as their relative positions; geometric of atoms that interact with them also have complex influences on $\mathbf{h}_{ij}$. Under thermal motions and rotational disturbances, the high-dimensional $\mathbf{h}_{ij}$ may vary in a vast space, which is difficult to be covered with a finite set of discrete training samples in the learning process. A potential alternative approach is to consider operators that maintain SO(3)-equivariance independent of the learning process from the perspective of group theory. However, as analyzed in Section 2, this approach introduces the limitations on the network’s expressive capabilities. This paper aims to resolve these challenges, achieving a harmonious integration of SO(3)-equivariance and non-linear expressive power in the deep Hamiltonian regression model.
@@
Figure 1: Left part: overview of the two-stage regression framework for Hamiltonian prediction, with the first stage extracting features with prior SO(3)-equivariance independent to the learning process and making baseline predictions, and the second stage enhancing accuracy via advanced non-linear mechanisms. The first-stage network aids the second-stage network in learning SO(3)-equivariance, while the second-stage network compensates for the expressiveness capability of first-stage network. Right part: The internal architecture of the non-linear 3D graph Transformer network.
4 Method
@@
As shown in Fig. 1, to harmonize SO(3)-equivariance and expressiveness for the deep learning paradigm, we propose a novel framework, namely HarmoSE, effectively hybridizing networks in two regression stages, from which the first-stage network, a group theory-informed network possessing SO(3)-equivariance prior to the learning process, provides essential foundations to the second stage in mastering SO(3)-equivariance, whereas the second-stage network, with highly expressive non-linear mappings, enriches the expressiveness capabilities of the whole framework. The combination of these two stages not only enhances expressiveness but also ensures robust equivariance to rotations of reference frames, bringing accurate predictions for electronic structure Hamiltonians despite rotational transformations.
4.1 Initial Features
@@
In our framework, the initial feature for the $i(1\leq i\leq N)$ th node is its node embedding, denoted as $\mathbf{z}_{i}$, a coordinate-independent SO(3)-invariant semantic embedding that marks its element type. Given the locality of the Hamiltonian [13], each atom $j$ in the local set $\Omega(i)$ within the cutoff radius of an atom $i$ form an edge with $i$. From each edge, a Hamiltonian is defined. The initial features for edge $(i,j)$ include both SO(3)-invariant encodings and SO(3)-equivariant encodings. The former includes edge embeddings $\mathbf{z}_{ij}$ marking the types of interacting atom pairs, as well as the distance features $\mathbf{d}_{ij}$ in the form of Gaussian functions [13]; the latter is spherical harmonics, denoted as $Y_{l}^{m}(\frac{\mathbf{r}_{ij}}{||\mathbf{r}_{ij}||})$, where $\frac{\mathbf{r}_{ij}}{||\mathbf{r}_{ij}||}$ describes the relative orientation between two atoms. Worth noting, a node and itself also possess a Hamiltonian. In other words, $i\in\Omega(i)$ and $\mathbf{e}_{ii}$ also forms a valid edge. In this case, $Y_{l}^{m}(\frac{\mathbf{r}_{ij}}{||\mathbf{r}_{ij}||})$ cannot be used to calculate directional features and a vector filled with 00 is used as a substitute. The above initial features are used to model the Hamiltonian which obeys SO(3)-equivariance, additionally, these features also inherently fulfill translational invariance with respect to global coordinates, another natural attribute of Hamiltonians.
4.2 The First Regression Stage
@@
In our framework, the primary role of the first regression stage is to extract node and edge representations with intrinsic SO(3)-equivariance independent to network parameters, and regress baseline Hamiltonian predictions, denoted as $\hat{\mathbf{h}}_{ij}$ in the direct sum state and $\hat{\mathbf{H}}_{ij}$ in the direct product state for each pair $(i,j)$, establishing a firm foundation on SO(3)-equivariance in both feature level and the regression target level. The encoding process of the $p$-th encoder group can be described by the following equation:
@@
where $\mathbf{n}^{(p)}_{i}$ and $\mathbf{e}^{(p)}_{ij}$ respectively denote the node and edge features from the $p$ th $(1\leq p\leq P)$ encoder group, $Gate(\cdot)$ is the gated activation function introduced in Section 2, $EquiLin(\cdot)$ denotes a combination of tensor operators consisting of linear scaling, element-wise sum, direct sum, direct product, as well as the Clebsch-Gordan decomposition, possessing parameter-independent SO(3)-equivariance guaranteed by group theory. These operators, serving as our first-stage backbone, have been comprehensively developed in previous works [20, 5, 6], achieving maturity in their equivariance capabilities, yet facing inherent limitations in non-linear expressiveness cannot be easily resolved within their mechanisms. Therefore, we focus more on the complementary mechanisms of the hybrid two-stage regression framework as well as the design of the second-stage network, aiming at rectifying the intrinsic weaknesses in non-linear expressiveness of the first-stage network while embodying robust SO(3)-equivariant capability.
4.3 The Second Regression Stage
@@
The second regression stage of our framework is designed to fully exploit non-linear mappings to enhance the expressive capability of the whole framework while capturing SO(3)-equivariance. For that purpose, as shown in the right part of Fig. 1, we propose a 3D graph Transformer which effectively models 3D atomic structures and predicts non-linear correction terms that complement the predictions from the first stage, achieving high-precision Hamiltonian prediction. Yet, one key problem to solve lies that, non-linear projections might not have parameter-independent guarantee on SO(3)-equivariance, forcing the second stage to capture equivariance through learning effective network parameters from the data. Thus, we must address the difficulty on empirical learning SO(3)-equivariance of Hamiltonians as discussed in Section 3. Our second-stage network adeptly resolves this issue, simultaneously capturing SO(3)-equivariance and enhancing the expressive capabilities. This is primarily attributed to three pivotal mechanisms we design.
@@
First, the second regression stage works in a cascaded manner, which means that its prediction target is not the entire Hamiltonian but a correction term $\Delta\hat{\mathbf{H}}_{ij}$, relative to the first stage’s output, i.e. the initial Hamiltonian estimate $\hat{\mathbf{H}}_{ij}$. The sum of these two stages’ outputs forms the final prediction of the Hamiltonian: $\mathbf{H}_{ij}=\hat{\mathbf{H}}_{ij}+\Delta\hat{\mathbf{H}}_{ij}$. Given that the predicted results of $\hat{\mathbf{H}}_{ij}$ are theoretically SO(3)-equivariant and numerically approximate to reasonable, the range of variations for the correction term $\Delta\hat{\mathbf{H}}_{ij}$, becomes smaller compared to $\hat{\mathbf{H}}_{ij}$. This reduces the complexity of the output space for the second stage and enhances the feasibility on implicitly capturing SO(3)-equivariance through data-driven learning for non-linear modules.

### v11

# Towards Harmonization of SO(3)-Equivariance and Expressiveness: a Hybrid Deep Learning Framework for Electronic-Structure Hamiltonian Prediction

@@ Abstract
Deep learning for predicting the electronic-structure Hamiltonian of quantum systems necessitates satisfying the covariance laws, among which achieving SO(3)-equivariance without sacrificing the non-linear expressive capability of networks remains unsolved. To navigate the harmonization between equivariance and expressiveness, we propose a deep learning method, i.e., HarmoSE, synergizing two distinct categories of neural mechanisms as a two-stage encoding and regression framework. The first stage corresponds to group theory-based neural mechanisms with inherent SO(3)-equivariant properties prior to the parameter learning process, while the second stage is characterized by a non-linear 3D graph Transformer network we propose, featuring high capability on non-linear expressiveness. The novel combination lies in the point that, the first stage predicts baseline Hamiltonians with abundant SO(3)-equivariant features extracted, assisting the second stage in empirical learning of equivariance; and in turn, the second stage refines the first stage’s output as a fine-grained prediction of Hamiltonians using powerful non-linear neural mappings, compensating for the intrinsic weakness on non-linear expressiveness capability of mechanisms in the first stage. Our method enables precise, generalizable predictions while capturing SO(3)-equivariance under rotational transformations, and achieves state-of-the-art performance in Hamiltonian prediction on six benchmark databases.
Introduction
@@
Deep learning methods (Schütt et al. 2019; Unke et al. 2021; Gu, Zhang, and Feng 2022; Li et al. 2022; Zhong et al. 2023; Gong et al. 2023) have emerged as a promising trend for predicting the electronic-structure Hamiltonian, an essential physical quantity in understanding a wide range of properties, including electronic-structures, magnetic properties, optics, transport, and numerous other properties. These methods have offered a way to bypass the computationally exhaustive self-consistent steps of the traditional Density Functional Theory (DFT) method (Hohenberg and Kohn 1964; Kohn and Sham 1965), thereby providing a viable pathway for the efficient simulation and design of large-scale atomic systems, laying the foundation for many down-stream applications (Zhang et al. 2023) in the information and energy areas.
@@
Despite these progresses, the Hamiltonian prediction task continues to present substantial challenges for deep learning techniques. High numerical accuracy is required to derive reasonable physical quantities, and furthermore, the fidelity of Hamiltonian predictions should not be confined to a specific coordinate system; rather, the results must demonstrate robust covariance and generalizability across various choices of reference frames. However, achieving 3D rotational equivariance, i.e., equivariance to the SO(3) group, is a tough target for a deep learning Hamiltonian prediction method. This difficulty arises because the Hamiltonian of each pair of atoms is usually high-dimensional, and its variation space under rotational disturbance is large. Consequently, it is difficult to cover the vast variability space they inhabit with rotations merely depending on parameter learning from discrete training samples. To address this, several works (Geiger and Smidt 2022; Gong et al. 2023) applied group theory-guided feature descriptors and tensor operators assuring inherent SO(3)-equivariance prior to the data-driven parameter learning process. Yet, to guarantee such SO(3)-equivariance independent to specific network parameters, these methods highly restricted the use of non-linear activation layers for SO(3)-equivariant features, leading to bottlenecks in expressiveness for complex non-linear mappings, limiting the accuracy achievable in predicting Hamiltonians. This dilemma, is also broadly prevalent in other 3D machine learning tasks where SO(3)-equivariance is highlighted, as analyzed by Zitnick et al. (2022).
@@
To harmonize SO(3)-equivariance and expressiveness for the prediction of electronic-structure Hamiltonians, this paper proposes a two-stage encoding and regression framework, i.e., HarmoSE, which combines mechanisms boasting parameter-independent prior SO(3)-equivariance with mechanisms featuring flexibility in non-linear expressiveness, overcoming the respective challenges of each categories of mechanisms, i.e., the limited non-linear expressive capability for the former as well as the difficulty of learning SO(3)-equivariance from data for the latter, through effective complementary strategies. Specifically, the first stage corresponds to the neural mechanisms constructed based on group theory with prior equivariant properties of 3D atomic systems, predicting an approximate value of the Hamiltonian, with abundant SO(3)-equivariant features provided. In the second stage, a highly expressive graph Transformer network we design, with no restrictions on non-linear activations, takes over. This network dynamically learns the 3D structural patterns of the atomic systems, compensates for the expressiveness shortcomings of the first-stage network arising from limited non-linear mappings, and refines the Hamiltonian values predicted in the first stage to enhance accuracy. Although this stage might not possess a parameter-independent prior SO(3)-equivariance due to its non-linearity, it is capable of capturing SO(3)-equivariance through learning effective network parameters with the help of three pivotal mechanisms. First, instead of directly regressing the entire Hamiltonians, the second stage aims to refine the Hamiltonian predictions from the first stage with corrective adjustments in a cascaded manner. The scopes of adjustments are smaller, lowing down the difficulties on non-linear learning of SO(3)-equivariance. Second, the second-stage network incorporates covariant features, including SO(3)-equivariant features extracted by the first-stage network and SO(3)-invariant features engineered by geometric knowledge, with its inputs to assist in the implicit learning of SO(3)-equivariance. Third, as the core of Transformer, the attention mechanism has potentials to adapt to geometric condition variations such as coordinate transformations, through its dynamic weighting strategy. Collectively, the combination of the two categories of neural mechanisms in the two stages allows the framework to overcome the challenges of each individual mechanism and make precise, generalizable, and SO(3)-equivariant predictions, being much more effective than simply increasing the parameter count for the network from one stage and fine-tuning it alone. Our method achieves state-of-the-art (SOTA) performance on Hamiltonian regression in six benchmark databases, demonstrating the superiority of our method. Particularly, our SOTA performance in the twisted samples, which exhibit both SO(3)-equivariance effects and variations in van der Waals (vdW) interactions due to the inter-layer rotations, comprehensively confirms the robust capability of our model in capturing the intrinsic SO(3)-equivariance of Hamiltonians as well as its strong non-linear expressive power to generalize to complex and dynamic 3D geometric structures of atomic systems.
Related Work
@@
As representative researches on equivariance to discrete rotational group, Dieleman et al. (Dieleman, Fauw, and Kavukcuoglu 2016) introduced cyclic symmetry operations into CNNs to achieve rotational equivariance; Ravanbakhsh et al. (Ravanbakhsh, Schneider, and Póczos 2017) explored parameter-sharing techniques for equivariance to discrete rotations; Kondor et al. (Kondor et al. 2018) developed equivariant representations via compositional methods and tensor theory; Zitnick et al. (Zitnick et al. 2022), Passaro et al. (Passaro and Zitnick 2023) and Liao et al. (Liao et al. 2024) introduced spherical harmonic bases for atomic modeling, focusing on rotational equivariance but limited to discrete sub-groups of SO(3) due to their sampling strategy. These approaches were very effective on discrete symmetries but sub-optimal on handling continuous 3D rotations. Focusing on equivariance to continuous rotational group, Jaderberg et al. (Jaderberg et al. 2015) and Cohen et al. (Cohen and Welling 2017) achieved considerable success in 2D image recognition tasks by modeling equivariance to in-plane rotations. However, their applications were limited within the scope of 2D tasks and did not fit for the more complex demands of equivariance to 3D continuous rotational group, i.e. SO(3), required in the Hamiltonian prediction task.
@@
In the field of researches on equivariance to SO(3), approaches like DeepH (Li et al. 2022) explored equivariance via a local coordinate strategy, which made inference within the fixed local coordinate systems built with neighboring atoms, then transferred the output according to equivariance rules to the corresponding global coordinates. However, due to a lack of in-depth exploration of SO(3)-equivariance at the neural mechanism level, this method faced challenges when the local coordinate system underwent rotational disturbances from non-rigid deformation, e.g. the inter-layer twist of bilayer structures. In contrast, methods like TFN (Thomas et al. 2018), SE(3)-Transformer (Fuchs et al. 2020), E3NN (Geiger and Smidt 2022), Equiformer (Liao and Smidt 2023), and DeephE3 (Gong et al. 2023), considered SO(3)-equivariance from the perspective of intrinsic mechanisms of neural networks, developed group theory-informed equivariant operations, such as linear scaling, element-wise sum, direct sum, direct product, and the Clebsch-Gordan decomposition, effectively applying in multiple atomic modeling tasks, where DeepHE3 (Gong et al. 2023) stood as a SOTA method across diverse atomic systems on Hamiltonian prediction. However, a common challenge across these methods lies in the fact that, to achieve inherent equivariance prior to the parameter learning process, they forbade the use of complex non-linear mappings like $Sigmoid$, $SiLU$, and $Softmax$ for SO(3)-equivariant features, significantly limiting the network’s expressive potential and causing a bottleneck in generalization performance. Although these approaches tried to enhance expressiveness via a gated activation function, where SO(3)-invariant features undergone through non-linear activation layers were used as gating coefficients that were multiplied with SO(3)-equivariant features, this mechanism, viewed from the perspective of equivariant features, amounted to a linear operation and did not fundamentally improve their expressive capability. For these methods, this equivariance-expressiveness dilemma remains an unsolved problem.
Preliminary
@@
In the study of symmetry on mathematical structures, An operation $A$ is equivariant with respect to $B$ if applying $B$ before or after $A$ has the same effect, expressed as: $A(B(x))=B(A(x))$. The key equivariance properties of Hamiltonians are the 3D rotational equivariance with respect to reference frame. Specifically, when the reference frame rotates by a rotation matrix denoted as $\mathbf{R}$, the edge of an atom pairs $(i,j)$ transforms from $\mathbf{r}_{ij}$ to $\mathbf{R}\cdot\mathbf{r}_{ij}$, and the Hamiltonians in the direct sum state transforms equivariantly from $\mathbf{h}_{ij}$ to $D(\mathbf{R})\cdot\mathbf{h}_{ij}$, where $D(\mathbf{R})$ is the Wigner-D matrix  [^1] Here we present SO(3)-equivariance under the direct sum state due to its simple vector form. For the equivalent formulation under the matrix-formed direct product state, please refer to Gong et al. (2023). The requirements for the fitting and generalization capability of a neural network $f_{nn}(\cdot)$ for Hamiltonian prediction can be formally expressed as: $f_{nn}(\{\mathbf{r}_{ij}|i\in Nodes,ij\in Edges\})\cong\{\mathbf{h}_{ij}\}$; moreover, the requirement on SO(3)-equivariance can be represented as: $f_{nn}(\{\mathbf{R}\cdot\mathbf{r}_{ij}|i\in Nodes,ij\in Edges\})\cong\{D(\mathbf{R})\cdot\mathbf{h}_{ij}\}$. It is crucial that $f_{nn}(\cdot)$ intrinsically captures SO(3)-equivariance to effectively generalize under rotational reference frames, and meanwhile, $f_{nn}(\cdot)$ must also possess sufficient expressive power to generalize across different types and structures of atomic systems and make accurate predictions.
@@
Figure 1: Left part: overview of the two-stage encoding and regression framework for Hamiltonian prediction. Right part: The internal architecture of the non-linear 3D graph Transformer network.
Method
@@
As shown in Fig. 1, to harmonize SO(3)-equivariance and expressiveness for Hamiltonian prediction, we propose a hybrid framework, i.e., i.e., HarmoSE, with two encoding and regression stages, from which the first-stage network, a group theory-informed network possessing SO(3)-equivariance prior to the learning process, provides essential foundations to the second stage in mastering SO(3)-equivariance, whereas the second-stage network, with highly expressive non-linear mappings, enriches the expressiveness capabilities of the whole framework. The combination of these two stages not only enhances expressiveness but also ensures robust equivariance to rotations of reference frames, bringing accurate predictions for electronic-structure Hamiltonians despite rotational transformations.
Initial Features
@@
In our framework, the initial feature for the $i(1\leq i\leq N)$ th node is its node embedding, denoted as $\mathbf{z}_{i}$, a coordinate-independent SO(3)-invariant semantic embedding that marks its element type. Given the locality of the Hamiltonian (Li et al. 2022), each atom $j$ in the local set $\Omega(i)$ within the cutoff radius of an atom $i$ form an edge with $i$. From each edge, a Hamiltonian is defined. The initial features for edge $(i,j)$ include both SO(3)-invariant encodings and SO(3)-equivariant encodings. The former includes edge embeddings $\mathbf{z}_{ij}$ marking the types of interacting atom pairs, as well as the distance features $\mathbf{d}_{ij}$ in the form of Gaussian functions (Li et al. 2022); the latter is spherical harmonics, denoted as $Y_{l}^{m}(\frac{\mathbf{r}_{ij}}{||\mathbf{r}_{ij}||})$ (Schrödinger 1926), where $\frac{\mathbf{r}_{ij}}{||\mathbf{r}_{ij}||}$ describes the relative orientation between two atoms.
The First Regression Stage
@@
In our framework, the primary role of the first encoding and regression stage is to extract node and edge representations with intrinsic SO(3)-equivariance independent to network parameters, and regress baseline Hamiltonian predictions, denoted as $\hat{\mathbf{h}}_{ij}$ in the direct sum state and $\hat{\mathbf{H}}_{ij}$ in the direct product state for each pair $(i,j)$, establishing a firm foundation on SO(3)-equivariance in both feature level and the regression target level. The encoding process of the $p$-th encoder group can be described by the following equation:
@@
where $\mathbf{n}^{(p)}_{i}$ and $\mathbf{e}^{(p)}_{ij}$ respectively denote the node and edge features from the $p$ th $(1\leq p\leq P)$ encoder group, $Gate(\cdot)$ is the gated activation function introduced in the Related Work Section, $EquiLin(\cdot)$ denotes a combination of tensor operators consisting of linear scaling, element-wise sum, direct sum, direct product, as well as the Clebsch-Gordan decomposition, possessing parameter-independent SO(3)-equivariance guaranteed by group theory. These operators, serving as our first-stage backbone, have been comprehensively developed in previous works (Thomas et al. 2018; Geiger and Smidt 2022; Gong et al. 2023), achieving maturity in their equivariance capabilities, yet facing inherent limitations in non-linear expressiveness cannot be easily resolved within their mechanisms. Therefore, we focus more on the complementary mechanisms of the hybrid two-stage encoding and regression framework as well as the design of the second-stage network, aiming at rectifying the intrinsic weaknesses in non-linear expressiveness of the first-stage network while embodying robust SO(3)-equivariant capability.
The Second Regression Stage
@@
The second encoding and regression stage of our framework is designed to fully exploit non-linear mappings to enhance the expressive capability of the whole framework while capturing SO(3)-equivariance. For that purpose, as shown in the right part of Fig. 1, we propose a 3D graph Transformer which effectively models 3D atomic structures and predicts non-linear correction terms that complement the predictions from the first stage, achieving high-precision Hamiltonian prediction. Yet, one key problem to solve is, non-linear projections might not have parameter-independent guarantee on SO(3)-equivariance, forcing the second stage to capture equivariance through learning effective network parameters from the data. Thus, the difficulty on empirical learning SO(3)-equivariance of Hamiltonians must be addressed. Our second-stage network adeptly resolves this issue, simultaneously capturing SO(3)-equivariance and enhancing the expressive capabilities. This is primarily attributed to three pivotal mechanisms we design.
@@
First, the second stage works in a cascaded manner for regression, which means that its prediction target is not the entire Hamiltonian but a correction term $\Delta\hat{\mathbf{H}}_{ij}$, relative to the first stage’s output, i.e. the initial Hamiltonian estimate $\hat{\mathbf{H}}_{ij}$. The sum of these two stages’ outputs forms the final prediction of the Hamiltonian: $\mathbf{H}_{ij}=\hat{\mathbf{H}}_{ij}+\Delta\hat{\mathbf{H}}_{ij}$. Given that the predicted results of $\hat{\mathbf{H}}_{ij}$ are theoretically SO(3)-equivariant and numerically approximate to reasonable, the range of variations for the correction term $\Delta\hat{\mathbf{H}}_{ij}$, becomes smaller compared to $\hat{\mathbf{H}}_{ij}$. This reduces the complexity of the output space for the second stage and enhances the feasibility on implicitly mastering SO(3)-equivariance through data-driven learning for non-linear modules.